OpenStudy (anonymous):

What is the solution to the equation 21 ^x – 3 = 14 ?

OpenStudy (anonymous):

i would double check that

OpenStudy (anonymous):

21^x-3=14 sum both sides 3 21^x=17 Put ln both sides(natural logarythm) ln21^x=17 Put down the exponent. xln21=17 divide both sides by ln21 x=ln17/ln21 Is ok now?

OpenStudy (anonymous):

is the problem \[21^x-3 =14 \] or\[21^(x-3)=14\] big difference

OpenStudy (anonymous):

im sory i ment \[21^{x-3}\]

OpenStudy (anonymous):

if it is the later you are in correct if its the first then you are correct

OpenStudy (anonymous):

answer is ln17/ln21..would u like the explainations